# Thermodynamic Entropic Uncertainty Relation

### v3

@@ Abstract
where $f\in\mathbb{R}$ is a constant force, $\xi(t)$ is the white Gaussian noise with the correlation $\braket{\xi(t)\xi(t^{\prime})}=\delta(t-t^{\prime})$, and $A$ is the noise intensity. We define the accumulated evidence as $y(\tau)=\int_{0}^{\tau}\partial_{t}x(t)dt$ and the observable is $\Phi(\Gamma)=1$, $-1$, and 00 when $y(\tau)\geq\theta$, $y(\tau)\leq-\theta$, and $-\theta<y(\tau)<\theta$, respectively. Here, our interest is in the relationship between the entropy of the selected decision $H[P(\Phi)]$ and the entropy production of the Langevin equation. Since $\Phi(\Gamma)$ satisfies the time-reversal antisymmetric condition [Eq. (7)], it satisfies Eqs. (14) and (15), implying that greater entropy production is necessary for lower decision entropy $H[P(\Phi)]$ under the steady-state condition. In decision making, it is well known that there exists a speed-accuracy tradeoff [32, 33], responding faster often reduces accuracy, while achieving higher accuracy lengthens decision time. Through the thermodynamic entropic uncertainty relation, we show quantitatively that another tradeoff, an energy-accuracy tradeoff, holds in the decision model.
@@
We perform numerical calculations for the diffusion decision model [Eq. (17)] to confirm Eq. (15) [26]. We calculate quantities, such as $\Sigma$ and $H[P(\Phi)]$, as a function of $f$ in Fig. 3 for different threshold settings (a) $\theta=0$ and (b) $\theta=0.5$. We consider the steady-state condition and calculate the probability distribution $P(\Phi)$ and $\Sigma$. In Fig. 3, $\Sigma+H[P(\Phi)]$ is represented by the solid line, where the dotted line indicates (a) $\ln 2$ or (b) $(1-P(0))\ln 2$. As shown in Figs. 3(a) and (b), $\Sigma+H[P(\Phi)]$ is always above $\ln 2$ or $(1-P(0))\ln 2$, respectively, confirming the validity of Eq. (15). It is clear that in the case of $\theta=0$, the values of $\Sigma+H[P(\Phi)]$ and the lower bound are closer than in the case of $\theta=0.5$. To examine how each of these components, $\Sigma$ and $H[P(\Phi)]$, behave as a function of $f$, $\Sigma$ and $H[P(\Phi)]$ are represented by the dashed line and the dot-dashed line, respectively. Regarding entropy production $\Sigma$, it falls below $\ln 2$ in the region where $f$ ranges from $0.0$ to $0.25$ in Fig. 3(a). On the other hand, the entropy of the observable $H[P(\Phi)]$ is decreasing as a function of $f$, and except for $f=0.0$, it is below $\ln 2$. From these results, we can conclude that the inequality $\Sigma+H[P(\Phi)]\geq\ln 2$ holds only when considering the sum, $\Sigma+H[P(\Phi)]$, that is, $\Sigma$ or $H[P(\Phi)]$ alone falls below $\ln 2$.
@@
Conclusion.—In this study, we established the thermodynamic entropic uncertainty relation that links the entropy production and the Shannon entropy of the observable. Our findings extend conventional thermodynamic uncertainty relations by incorporating measures based on entropy, highlighting the role of Shannon entropy of observables in stochastic thermodynamics. The derived inequality formalizes a fundamental trade-off between entropy production and the asymmetry of the observable distribution. Through the application to the diffusion decision model, we revealed that accurate decision necessarily require higher entropy production. This framework provides a deeper understanding of nonequilibrium thermodynamics and expands the application of entropy-based uncertainty relations in stochastic systems. We have considered the Kullback-Leibler divergence with respect to trajectories $\Gamma$, but it is also possible to start from the divergence between different quantities (see the supplementary material [26]). Another direction of expansion is towards quantum systems. In recent years, the thermodynamic uncertainty relations in quantum systems [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47] have garnered significant attention. In particular, the thermodynamic uncertainty relations within the framework of continuous measurement are closely related to those in classical stochastic processes. This direction presents future challenges.
Acknowledgements.
This work was supported by JSPS KAKENHI Grant Number JP23K24915.
Appendix A Derivation of the main result [Eq. (14)]
@@ Acknowledgements.
Let us introduce the absolute random variable of $\Phi$, which is expressed by $|\Phi|$. The probability distribution of $|\Phi|$ is given by
@@
Using the monotonicity of Kullback-Leibler divergence [Eq. (5)] and the time-reversal property of $\Phi$ [Eq. (7)], we have
@@
Here, the cross entropy is evaluated as $C[P(\Phi),P(-\Phi)]=-\sum_{\phi}P(\Phi=\phi)\ln P(\Phi=-\phi)$. Since the cross entropy term is no less than the entropy, it is non-negative, $C[P(\Phi),P(-\Phi)]\geq H[P(\Phi)]\geq 0$. An important observation is that $C[P(\Phi),P(-\Phi)]$ is even bounded from below by a positive term. We calculate $H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$ as follows:
@@
Let us consider the function $f(x)=\ln\frac{x}{1+x}$. Since $f(x)$ is concave for $x>0$, by using the Jensen inequality, the following relation holds:
@@
$P(|\Phi|=\phi)=P(\Phi=\phi)+P(\Phi=-\phi).$ (25)
We do not consider $P(|\Phi|=0)$, because such measure is 00 for the smooth probability density. When $P(\Phi)$ includes the contribution of the delta function in $\Phi=0$, this is not the case. Following the same procedure as the discrete case, we obtain
@@
Appendix B Symmetry entropy and Jensen-Shannon divergence
The Jensen-Shannon divergence is defined by
$\mathrm{JS}[P(X)\|P(Y)]$ $\equiv\frac{1}{2}D\left[P(X)\|\frac{P(X)+P(Y)}{2}\right]$
$+\frac{1}{2}D\left[P(Y)\|\frac{P(X)+P(Y)}{2}\right],$ (27)
@@ Appendix B Symmetry entropy and Jensen-Shannon divergence
Appendix C Proof of Eq. (9)
In this section, we prove Eq. (9). Here, we show the relation for $P(\Phi=0)\geq 0$. The relation which we want to show is given by
$0\leq\Lambda[P(\Phi)]\leq[1-P(\Phi=0)]\ln 2$ (30)
The first inequality part corresponds to $H[P(|\Phi|)]\leq H[P(\Phi)]$. For random variables $X$ and $Y$, the following relation holds:
$H[P(X),P(Y)]$ $=H[P(X)|P(Y)]+H[P(Y)]$
$=H[P(Y)|P(X)]+H[P(X)],$ (31)
where $H[P(Y)|P(X)]$ is the conditional entropy:
$H[P(Y)|P(X)]$
$\equiv-\sum_{x,y}P(X=x,Y=y)\ln P(Y=y|X=x).$ (32)
Substituting $X=\Phi$ and $Y=|\Phi|$, we obtain
$H[P(|\Phi|)|P(\Phi)]+H[P(\Phi)]$ $=H[P(\Phi)]$
$=H[P(\Phi)|P(|\Phi|)]+H[P(|\Phi|)]$
$\geq H[P(|\Phi|)],$ (33)
where we used $H[P(|\Phi|)|P(\Phi)]=0$. Equation (33) proves the first inequality part of Eq. (30). Note that, for the continuous case, this proof does not work, as the continuous conditional entropy $H[P(\Phi)|P(|\Phi|)]$ may take negative values. However, we can show that the first inequality part also holds for the continuous case by a direct calculation. Specifically, $H[P(\Phi)]-H[P(|\Phi|)]$ is
$H[P(\Phi)]-H[P(|\Phi|)]$
[…]
$=\int_{0}^{\infty}d\phi\,\Bigl{[}[P(\phi)+P(-\phi)]\ln\left[P(\phi)+P(-\phi)\right]$
$-P(\phi)\ln P(\phi)-P(-\phi)\ln P(-\phi)\Bigr{]}.$ (34)
Because $(a+b)\ln(a+b)-a\ln a-b\ln b>0$ for $a>0$ and $b>0$, $H[P(\Phi)]-H[P(|\Phi|)]\geq 0$ is proved.
$H[P(|\Phi|)]-H[P(\Phi)]$
$\geq-\ln 2+P(0)\ln 2,$ (35)
where we again used the Jensen inequality. Equation (35) proves the second inequality part of Eq. (30).

### v5

@@ IV Example
where $f\in\mathbb{R}$ is a constant force, $\xi(t)$ is the white Gaussian noise with the correlation $\braket{\xi(t)\xi(t^{\prime})}=\delta(t-t^{\prime})$, and $A$ is the noise intensity. We define the accumulated evidence as $y(\tau)=\int_{0}^{\tau}\partial_{t}x(t)dt$ and the observable is
$\Phi(\Gamma)=\begin{cases}+1\;\;(\text{yes})&{\displaystyle y(\tau)\geq\theta}\\ -1\;\;(\text{no})&{\displaystyle y(\tau)\leq-\theta}\\ 0\;\;(\text{neutral})&\mathrm{otherwise}\end{cases}.$ (18)
Explicit expressions of the probability distribution of $y$ are shown in Appendix E. The entropy production during $0,\tau]$ is given by
$\Sigma=\frac{\tau f^{2}}{A}.$ (19)
The thermodynamic entropic uncertainty relation loses its predictive power when the entropy production $\Sigma$ exceeds the threshold $\ln 2$. For the diffusion decision model, this occurs when the time parameter $\tau$ exceeds the critical value $\tau^{*}$ defined as
$\tau^{*}\equiv\frac{A}{f^{2}}\ln 2.$ (20)
@@
Here, our interest is in the relationship between the entropy of the selected decision $H[P(\Phi)]$ and the entropy production $\Sigma$ of the Langevin equation. Since $\Phi(\Gamma)$ satisfies the time-reversal antisymmetric condition [Eq. (7)], it satisfies Eqs. (14) and (15), implying that greater entropy production is necessary for lower decision entropy $H[P(\Phi)]$ under the steady-state condition. In decision making, it is well known that there exists a speed-accuracy trade-off [33, 34], responding faster often reduces accuracy, while achieving higher accuracy lengthens decision time. Through the thermodynamic entropic uncertainty relation, we show quantitatively that another trade-off, an energy-accuracy trade-off, holds in the decision model.
@@
We perform numerical calculations for the diffusion decision model [Eq. (17)] to confirm Eq. (15) [26]. We calculate quantities, such as $\Sigma$ and $H[P(\Phi)]$, as a function of $f$ in Figs. 4(a) and (b) for different threshold settings (a) $\theta=0$ and (b) $\theta=0.5$. We consider the steady-state condition. In Figs. 4(a) and (b), $\Sigma+H[P(\Phi)]$ is represented by the solid line, where the dotted line indicates (a) $\ln 2$ or (b) $(1-P(0))\ln 2$. As shown in Figs. 4(a) and (b), $\Sigma+H[P(\Phi)]$ is always above $\ln 2$ or $(1-P(0))\ln 2$, respectively, confirming the validity of Eq. (15). It is clear that in the case of $\theta=0$, the values of $\Sigma+H[P(\Phi)]$ and the lower bound are closer than in the case of $\theta=0.5$. To examine how each of these components, $\Sigma$ and $H[P(\Phi)]$, behave as a function of $f$, $\Sigma$ and $H[P(\Phi)]$ are represented by the dashed line and the dot-dashed line, respectively. Regarding entropy production $\Sigma$, it falls below $\ln 2$ in the region where $f$ ranges from $0.0$ to $0.25$ in Fig. 4(a). On the other hand, the entropy of the observable $H[P(\Phi)]$ is decreasing as a function of $f$, and except for $f=0.0$, it is below $\ln 2$. From these results, we can conclude that the inequality $\Sigma+H[P(\Phi)]\geq\ln 2$ holds only when considering the sum, $\Sigma+H[P(\Phi)]$, that is, $\Sigma$ or $H[P(\Phi)]$ alone falls below $\ln 2$.
We study the $\tau$ dependence of the entropic quantities. We calculate the entropic quantities, such as $\Sigma$ and $H[P(\Phi)]$, as a function of $\tau$ in Fig. 4(c). The meaning of each line is the same as in Fig. 4(a). Since the entropy production increases linearly as a function of $\tau$ [Eq. (19)], we see that the predictive power of the thermodynamic entropic uncertainty relation is lost around $\tau^{*}=2.8$ [Eq. (20)] for this parameter setting.
Figure 5: Results for the diffusion decision model with the continuous-output observable [Eq. (21)]. Panels (a) and (b) show the entropic quantities versus $f$ for $A=1$ and $A=0.1$, respectively, with $\tau$ fixed at $10$. The blue dashed line shows the entropy production $\Sigma$, and the purple dot-dashed line represents $\Lambda[P(\Phi)]$, the symmetry entropy. Their sum, $\Sigma+\Lambda[P(\Phi)]$, is also depicted with the red solid line. The orange dotted line displays $\ln 2$, which serves as the lower bound for $\Sigma+\Lambda[P(\Phi)]$ according to Eq. (14).
Next, we consider the continuous-output observable case, where the observable is
$\Phi(\Gamma)=y(\tau).$ (21)
Equation (21) obviously satisfies the time-reversal antisymmetry condition. We numerically calculate the entropic quantities. We show $\Sigma$ and $\Lambda[P(\Phi)]$ as a function of $f$ in Fig. 5 for different noise intensity settings (a) $A=1$ and (b) $A=0.1$. Figure 5 shows the entropic quantities: $\Sigma+\Lambda[P(\Phi)]$ (solid line), $\ln 2$ (dotted line), $\Lambda[P(\Phi)]$ (dot-dashed line), and $\Sigma$ (dashed line). Again, Figs. 5(a) and (b) demonstrate that the sum $\Sigma+\Lambda[P(\Phi)]$ consistently exceeds $\ln 2$, supporting the correctness of Eq. (14). Specifically, when the noise intensity $A$ is set to the smaller value, as shown in Fig. 5(b), the symmetry entropy $\Lambda[P(\Phi)]$ approaches zero as $f$ increases. When $A$ is smaller, the entropy production $\Sigma$ increases, leading to a rapid increase in the value of $\Sigma+\Lambda[P(\Phi)]$ as a function of $f$. In Fig. 5(a), in the range where $f$ varies from $0.0$ to $0.25$, the entropy production $\Sigma$ is less than $\ln 2$. The symmetry entropy $\Lambda[P(\Phi)]$ decreases with increasing $f$ and remains below $\ln 2$ for all values except when $f=0.0$. Based on these observations, it is evident that the inequality $\Sigma+\Lambda[P(\Phi)]\geq\ln 2$ is satisfied only when considering the combined terms, $\Sigma+\Lambda[P(\Phi)]$. Individually, both $\Sigma$ and $\Lambda[P(\Phi)]$ drop below $\ln 2$.
V Conclusion
@@
In this study, we established the thermodynamic entropic uncertainty relation that links the entropy production and the Shannon entropy of the observable. Our findings extend conventional thermodynamic uncertainty relations by incorporating measures based on entropy, highlighting the role of Shannon entropy of observables in stochastic thermodynamics. The derived inequality formalizes a fundamental trade-off between entropy production and the asymmetry of the observable distribution. Through the application to the diffusion decision model, we revealed that accurate decisions necessarily require higher entropy production. This framework provides a deeper understanding of nonequilibrium thermodynamics and expands the application of entropy-based uncertainty relations in stochastic systems. We have considered the Kullback-Leibler divergence with respect to trajectories $\Gamma$, but it is also possible to start from the divergence between different quantities. Another direction of expansion is towards quantum systems. In recent years, the thermodynamic uncertainty relations in quantum systems [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48] have garnered significant attention. In particular, the thermodynamic uncertainty relations within the framework of continuous measurement are closely related to those in classical stochastic processes. This direction presents future challenges.
Appendix A Derivation of the main result [Eqs. (10) and (14)]
@@
Let us introduce the absolute value of the random variable $\Phi$, denoted by $|\Phi|$. The probability distribution of $|\Phi|$ is given by
@@
Using the monotonicity of the Kullback-Leibler divergence [Eq. (5)] and the time-reversal property of $\Phi$ [Eq. (7)], we have
@@
Here, the cross entropy is evaluated as $C[P(\Phi),P(-\Phi)]=-\sum_{\phi}P(\Phi=\phi)\ln P(\Phi=-\phi)$. Since the cross entropy term is no less than the entropy, it is non-negative, $C[P(\Phi),P(-\Phi)]\geq H[P(\Phi)]\geq 0$. The important observation is that $C[P(\Phi),P(-\Phi)]$ is even bounded from below by a positive term. We calculate $H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$ as follows:
@@
Let us consider the function
$\mathfrak{f}(x)=\ln\frac{x}{1+x}.$ (26)
Since $\mathfrak{f}(x)$ is concave for $x>0$, by using the Jensen inequality, the following relation holds:
@@
$P(|\Phi|=\phi)=P(\Phi=\phi)+P(\Phi=-\phi)\;\;(\phi\geq 0).$ (30)
We do not consider $P(|\Phi|=0)$, because this set has measure zero for the smooth probability density. When $P(\Phi)$ includes the contribution of the delta function in $\Phi=0$, this is not the case. Following the same procedure as the discrete case, we obtain
@@
Appendix B Proof of Eqs. (9) and (13)
In this section, we prove Eqs. (9) and (13). Here, we show the relation for $P(\Phi=0)\geq 0$. The relation that we want to show is given by
$0\leq\Lambda[P(\Phi)]\leq[1-P(\Phi=0)]\ln 2.$ (32)
The first inequality part corresponds to $H[P(|\Phi|)]\leq H[P(\Phi)]$. Specifically, $H[P(\Phi)]-H[P(|\Phi|)]$ is
$H[P(\Phi)]-H[P(|\Phi|)]$
$=\sum_{{\phi>0}}[P(\phi)+P(-\phi)]\ln\left[P(\phi)+P(-\phi)\right]$
${+P(0)\ln P(0)}-\sum_{\phi}P(\phi)\ln P(\phi)$
$=\sum_{\phi>0}\Bigl[[P(\phi)+P(-\phi)]\ln\left[P(\phi)+P(-\phi)\right]$
$-P(\phi)\ln P(\phi)-P(-\phi)\ln P(-\phi)\Bigr]\geq 0,$ (33)
where we used $(a+b)\ln(a+b)-a\ln a-b\ln b>0$ for $a>0$ and $b>0$ in the last line. $(a+b)\ln(a+b)-a\ln a-b\ln b=0$ if and only if $a=0$ or $b=0$. This indicates that Eq. (33) saturates if and only if $P(\phi)=0$ or $P(-\phi)=0$ for all $\phi$ except $P(0)$.
$H[P(|\Phi|)]-H[P(\Phi)]$
$\geq-\ln 2+P(0)\ln 2,$ (34)
where we again used the Jensen inequality. Equation (34) proves the second inequality part of Eq. (32). Equality in Eq. (34) holds if and only if $P(-\phi)/P(\phi)=1$.
Appendix C Symmetry entropy and Jensen-Shannon divergence
We derive Eq. (11). Although we assume $P(\Phi=0)=0$ in Eq. (11), the derivation below also applies to the general case, including $P(\Phi=0)>0$. The Jensen-Shannon divergence is defined by
$\mathrm{JS}[P(X)\|P(Y)]$ $\equiv\frac{1}{2}D\left[P(X)\middle\|\frac{P(X)+P(Y)}{2}\right]$
$+\frac{1}{2}D\left[P(Y)\middle\|\frac{P(X)+P(Y)}{2}\right],$ (35)
@@ Appendix C Symmetry entropy and Jensen-Shannon divergence
Appendix D Equality condition
We consider the equality condition of the thermodynamic entropic uncertainty relation [Eqs. (10) and (14)]. The derivation of the thermodynamic entropic uncertainty relation employs two key inequalities:
  1. 1.
    Monotonicity of the Kullback-Leibler divergence under coarse graining.
  2. 2.
    The Jensen inequality applied to a strictly concave function.
To saturate the thermodynamic entropic uncertainty relation, both inequalities must be saturated simultaneously. The distribution $P(\phi)$ is defined by
$P(\phi)=\sum_{\Gamma}\delta\bigl(\Phi(\Gamma)-\phi\bigr)\mathcal{P}(\Gamma).$ (38)
Then the monotonicity states
$D\bigl[\mathcal{P}(\Gamma)\|\mathcal{P}(\Gamma^{\dagger})\bigr]\geq D\bigl[P(\phi)\|P(-\phi)\bigr].$ (39)
Equality holds if and only if the log-ratio depends on $\Gamma$ solely through $\Phi(\Gamma)$, i.e., there exists a function $h$ such that
$\frac{\mathcal{P}(\Gamma)}{\mathcal{P}(\Gamma^{\dagger})}\;=\;h\!\bigl(\Phi(\Gamma)\bigr).$ (40)
There are two scenarios in which Eq. (40) is satisfied:
Case 1 (Equilibrium)
$\mathcal{P}(\Gamma)=\mathcal{P}(\Gamma^{\dagger})$ for all $\Gamma$.
Case 2 (Bijectivity)
$\Phi(\Gamma)$ is a bijective function.
Next, we consider the Jensen inequality part. We applied the Jensen inequality to the strictly concave function $\mathfrak{f}(x)$ defined in Eq. (26). The equality of the Jensen inequality is saturated if and only if the following condition is satisfied:
$\frac{P(\phi)}{P(-\phi)}\;=\;\text{constant}.$ (41)
Since Eq. (41) should hold for all $\phi$, Eq. (41) is equivalently stated as
$P(\phi)=P(-\phi).$ (42)
To saturate the thermodynamic entropic uncertainty relation, we require both (40) and (42). We now show that this forces equilibrium:
  • In Case 1, $\mathcal{P}(\Gamma)=\mathcal{P}(\Gamma^{\dagger})$ already defines the equilibrium. Moreover, equilibrium implies $P(\phi)=P(-\phi)$, so Eq. (42) is automatically satisfied.
  • In Case 2, $\Phi$ is bijective, so Eq. (40) allows any $\mathcal{P}(\Gamma)$ expressed via $\Phi(\Gamma)$. But Eq. (42) then demands $\mathcal{P}(\Gamma)/\mathcal{P}(\Gamma^{\dagger})$ be constant, again forcing $\mathcal{P}(\Gamma)=\mathcal{P}(\Gamma^{\dagger})$.
In summary, the thermodynamic entropic uncertainty relation [Eqs. (10) and (14)] is saturated if and only if the system is in equilibrium.
Appendix E Diffusion decision model
E.1 Model
In the main text, we consider the diffusion decision model, where the evidence accumulation process is modeled by Eq. (17). Let $P(x;t)$ be the probability density of $x$ at time $t$. The corresponding Fokker-Planck equation is
$\frac{\partial P(x;t)}{\partial t}=-\frac{\partial}{\partial x}[fP(x;t)]+A\frac{\partial^{2}P(x;t)}{\partial x^{2}}.$ (43)
The probability current of Eq. (43) is
$J(x;t)\equiv fP(x;t)-A\frac{\partial P(x;t)}{\partial x}.$ (44)
A trajectory $\Gamma$ corresponds to a realization of the Langevin dynamics within the time interval $0,\tau]$. In the main text, we define the observable $\Phi(\Gamma)$ for a decision-making process. The probability density of $y(t)$ is given by the Gaussian distribution [49]:
$P(y;t)=\frac{1}{\sqrt{4\pi At}}\exp\left[-\frac{\left(y-ft\right)^{2}}{4At}\right].$ (45)
E.2 Discrete case
In the discrete model, the observable is defined in Eq. (18). From Eq. (45), the observable probability distribution at time $\tau$ is
$P(\Phi=+1)$ $=\int_{\theta}^{\infty}P(y;t)dy=\frac{1}{2}\mathrm{erf}\left(\frac{ft-\theta}{2\sqrt{At}}\right)+\frac{1}{2},$ (46)
$P(\Phi=0)$ $=\int_{-\theta}^{\theta}P(y;t)dy=-\frac{1}{2}\mathrm{erf}\left(\frac{ft-\theta}{2\sqrt{At}}\right)$
$+\frac{1}{2}\mathrm{erf}\left(\frac{ft+\theta}{2\sqrt{At}}\right),$ (47)
$P(\Phi=-1)$ $=\int_{-\infty}^{-\theta}P(y;t)dy=\frac{1}{2}-\frac{1}{2}\mathrm{erf}\left(\frac{ft+\theta}{2\sqrt{At}}\right).$ (48)
The probabilities in Eqs. (46)–(48) are used to calculate $\Lambda[P(\Phi)]$.
E.3 Continuous case
The symmetry entropy is given by $\Lambda[P(\Phi)]=H[P(\Phi)]-H[P(|\Phi|)]$ [Eq. (8)]. For the observable given by Eq. (21), each of the terms in $\Lambda[P(\Phi)]$ is
$H[P(\Phi)]$ $=-\int_{-\infty}^{\infty}dy\,P(y;\tau)\ln P(y;\tau),$ (49)
$H[P(|\Phi|)]$ $=-\int_{0}^{\infty}dy\,\Bigl[\left[P(y;\tau)+P(-y;\tau)\right]$
$\times\ln\left[P(y;\tau)+P(-y;\tau)\right]\Bigr],$ (50)
where $P(y;t)$ is Eq. (45).
Acknowledgements.
This work was supported by JSPS KAKENHI Grant Number JP23K24915.
DATA AVAILABILITY
The data that support the findings of this article are openly available [50].